OpenStudy (anonymous):

5!-4!=[1] evaluate........bruh i NEVER SEEN THIS IN MY LIFE

OpenStudy (baby_bear69):

5! is 5x4x3x2x1 4! is 4x3x2x1 solve that first

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

150 and and 24 ok what i do next @Baby_Bear69

OpenStudy (baby_bear69):

5! is 120 Take 120-24

OpenStudy (anonymous):

120 ?? sorry dang

OpenStudy (baby_bear69):

Haha yep

OpenStudy (anonymous):

96

OpenStudy (baby_bear69):

so 120-24 = 96 therefore 5!-4! = 96 (:

OpenStudy (anonymous):

tank ewe sooo much @Baby_Bear69

OpenStudy (baby_bear69):

No problem!

OpenStudy (xapproachesinfinity):

! this symbol is called factorial i has many applications in mathematics the mostly widely known one is in counting

OpenStudy (empty):

I wouldn't do it this way, evaluating it is a waste of time! Use this property of factorials: \[n! = n* (n-1)!\] Then we can write it a little quicker now \[5!-4! = 5*4!-4!= (5-1)4! = 4*4*3*2*1 = 96\]